OpenStudy (anonymous):

2x-y=6 -3y=-6x+18 using subititution

OpenStudy (diyadiya):

2x-y=6 -- (1) -3y=-6x+18 --(2) from (1) 2x=y+6 ( by adding y on both sides) x=(y+6)/2 ( dividing both sides by 2 ) now substitute the value of x in (2)

OpenStudy (goformit100):